OpenStudy (anonymous):

The surface area of a cone whose radius is 5 cm and whose height is 12 cm is _____ cm². Use 3.14 for π. 264.2 266.9 282.6 296.4

jimthompson5910 (jim_thompson5910):

use the formula shown here http://0.tqn.com/d/math/1/0/t/L/conerr.jpg

jimthompson5910 (jim_thompson5910):

first you need s -------------- r^2 + h^2 = s^2 5^2 + 12^2 = s^2 25 + 144 = s^2 s^2 = 169 s = 13 ------------- now onto the surface area SA = pi*r^2 + pi*r*s SA = 3.14*5^2 + 3.14*5*13 SA = ???

OpenStudy (anonymous):

204.1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

SA = 3.14*5^2 + 3.14*5*13 SA = 282.6